How I make slides in L^{A}TEX

Dennis Chao

Department of Computer Science

University of New Mexico

dlchao@cs.unm.edu

Outline

August 2003 • Introduction

• Using the foils package • Conclusions

Introduction

Why I make slides in L^{A}TEX:

• Don’t like PowerPoint

• Can import text and figures from other L^{A}TEX documents • Looks ok • Introduction

• Using the foils package • Conclusions

L^{A}TEX packages

I think you just need the foils package, available at www.ctan.org. Also getting the pdftricks package is recommended.

Compiling the document

I put conditionals in the source so L^{A}TEX can be used, but I use pdflatex for the final output because:

• Acroread is available on all platforms, so pdfs are a portable presentation format. Acroread has a full-screen mode that looks good.

• PDFs are nice because web search engines can parse them. This means that you can put your talk slides on the web and not only can you view them in a web browser using the acroread plugin, they can be found by people searching the web.

• I often use lots of figures in my slides, so it is nice that pdflatex can read jpg files.

• Introduction

• Using the foils package • Conclusions

Conclusions

Making slides in L^{A}TEX is easy.